The sum of two numbers is 137. The difference between the numbers is 29. What are the two numbers?
Answer by adamchapman(301)

Call the 2 numbers "x" and "y":
x+y=137...................(1)
x-y=29....................(2)
rearrange (2) to give y in terms of x:
y=x-29....................(3)
substitute (3) into (1) for x:
x+(x-29)=137
2x-29=137
2x=166
x=83......................(4)
substitute (4) into (3) for x:
y=83-29
y=54
I hope this helps,
Adam
P.S. please visit my website, it may be helpful to you. The address is www.geocities.com/quibowibbler